# Ergodicity, KAM, FPUT

### v1

@@ 1 Ergodicity
$\delta x(t)=x^{\prime}(\frac{i^{\prime}}{i}t)-x(t){\buildrel def\over{=}}x^{\prime}(i^{\prime}\varphi)-x(i\varphi)\hfill$ (1.4)
@@
The variation $\delta({\overline{K}}+{\overline{V}}+{\overline{V}_{ext}}){\buildrel def\over{=}}dU+dL$, with $U{\buildrel def\over{=}}{\overline{K}}+{\overline{V}}$ and $dL{\buildrel def\over{=}}d{\overline{V}_{ext}}$, has a natural interpretation of heat $\delta Q$ in/out in the process $x\to x^{\prime}$.  [^2] Because $\delta({\overline{K}}+{\overline{V}}){\buildrel def\over{=}}dU$ and $\delta{\overline{V}_{ext}}{\buildrel def\over{=}}\delta L$ are the variation of the internal energy $U$ and $dL$ is the work that the system performs: i.e. $\delta Q=\delta U+\delta L$.
@@
${{}={\bf-}2\frac{\delta i}{i}{\overline{K}}+{2\delta_{\xi}{{\overline{K}}}}+{\mbox{\boldmath$\delta$}}_{\xi}{\bf({-{{\overline{K}}}}+{\overline{V}}+{\overline{V}_{ext}})}\equiv{\bf+}2\frac{\delta i}{i}{\overline{K}}+2\delta{{\overline{K}}}+{\bf 0}}$
@@
${\frac{\delta Q}{{\overline{K}}}}=+2\frac{\delta i}{i}+2\frac{\delta{\overline{K}}}{{\overline{K}}}={2\delta\log(i{\overline{K}})}{\buildrel def\over{=}}{\bf\delta S}$ (1.7)
@@ 2 KAM
$H_{0}({\bf A},{\mbox{\boldmath$\alpha$}})=\frac{1}{2}({\bf A}\cdot J_{0}{\bf A})+{\mbox{\boldmath$\omega$}}_{0}\cdot{{\bf A}}+f_{0}({\bf A},{\mbox{\boldmath$\alpha$}})\equiv h_{0}+f_{0}$ (2.1)
real analytic for $({\bf A},{\mbox{\boldmath$\alpha$}})\in({\mathcal{D}}_{\rho}\times{\mathcal{T}}^{N})$, $N\geq 2$, with: ${\mathcal{D}}_{\rho}=\{{\bf A}\in R^{N},|A_{j}|<\rho\}$, ${\mathcal{T}}^{N}$ = $N$-dimensional torus $0,2\pi]^{N}$, identified with unit polydisk, $\{{\bf z}|\,z_{j}=e^{i\alpha_{j}},j=1,\ldots,N\}$, ${\mbox{\boldmath$\omega$}}_{0}\in R^{N}$ and $J_{0}$ is a $N\times N$ matrix, positive for simplicity, with eigenvalues in $J_{0,-},J_{0,+}]$.
@@
${{\mathcal{C}}_{\rho_{0},\kappa_{0}}}$ ${{}{\buildrel def\over{=}}\{({\bf A},{\bf z})||A_{j}|\,\leq\,\rho_{0},\ e^{-\kappa_{0}}\leq|e^{i\alpha_{j}}|\leq e^{\kappa_{0}},j=1,\ldots,N\}\subset{\mathcal{C}}^{2N}}$
@@
${||f}$ ${{}||_{\rho_{0},\kappa_{0}}{\buildrel def\over{=}}\max_{{\mathcal{C}}_{\rho_{0},\kappa_{0}}}|f({\bf A},{\bf z})|,\qquad\forall f\ {\rm holomorphic\ in}\ {\mathcal{C}}_{\rho_{0},\kappa_{0}}}$
@@
The idea is to focus attention on the center of ${\mathcal{C}}_{\rho_{0},\kappa_{0}}$ where, if $\varepsilon_{0}=0$, a motion (“free motion”: $({\bf A},{\mbox{\boldmath$\alpha$}})\to({\bf A},{\mbox{\boldmath$\alpha$}}+{\mbox{\boldmath$\omega$}}_{0}t)$) takes place which is quasi periodic “with spectrum” ${\mbox{\boldmath$\omega$}}_{0}$. This is done by studying motions in a small polydisk ${\mathcal{C}}_{\widetilde{\rho},\widetilde{\kappa}}({\bf a})\subset{\mathcal{C}}_{\rho_{0},\kappa_{0}}$, eccentric if ${\bf a}\neq{\bf 0}$.
Thus motions starting in ${\mathcal{C}}(\widetilde{\rho},\widetilde{\kappa})({\bf a})\subset{\mathcal{C}}(\rho_{0},\kappa_{0})$ near the center of ${\mathcal{C}}(\rho_{0},\kappa_{0})$ can be studied as “through a microscope”: in the good cases (i.e. under suitable assumption on the initial parameters $J_{0},{\mbox{\boldmath$\omega$}}_{0}$ and $f_{0}$) the Hamiltonian will turn out to be expressible (after a further coordinate change, to turn ${\mathcal{C}}(\widetilde{\rho},\widetilde{\kappa})({\bf a})$ into ${\mathcal{C}}(\rho_{1},\kappa_{1})$, centered at 0) in a form substantially closer to that of a quasi periodic oscillator (described by its “normal” Hamiltonian ${\mbox{\boldmath$\omega$}}_{0}\cdot{\bf A}$, or ${\mbox{\boldmath$\omega$}}_{0}\cdot{\bf A}+\frac{1}{2}{{\bf A}}\cdot J_{0}{{\bf A}}$, in the variables ${\bf A},{\mbox{\boldmath$\alpha$}}$).
Iterating the process the Hamiltonian changes but, remaining analytic in a polydisk $C_{\rho_{n},\kappa_{n}}$, converges to that of a harmonic oscillator: the interpretation will be that, looking very carefully in the vicinity of the “unperturbed” torus ${\mathcal{T}}_{{\mbox{\boldmath$\omega$}}_{0}}=\{{\bf A}={\bf 0},{\mbox{\boldmath$\alpha$}}\in[0,2\pi]^{N}\}$, the perturbed Hamiltonian exhibits with increasing precision a harmonic motion with spectrum ${\mbox{\boldmath$\omega$}}_{0}$.
@@
The “spectrum” ${\mbox{\boldmath$\omega$}}_{0}$ will be supposed “diophantine”, i.e. for some $C>0$ it is, denoting $\mathcal{Z}^{N}$ the lattice of the integers, for all ${\bf 0}\neq{\mbox{\boldmath$\nu$}}\in\mathcal{Z}^{N}$:
$|{\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}|\,\geq\,C\,|{\mbox{\boldmath$\nu$}}|^{-N},\qquad|{\mbox{\boldmath$\nu$}}|\equiv\sum_{i=1}^{N}|\nu_{i}|>0$ (2.3)
The inequality Eq.2.3 will be repeatedly used to define canonical transformations $({{\bf A}},{\mbox{\boldmath$\alpha$}})\,{\kern-1.0pt\leftarrow\kern-5.0pt\to\kern-1.0pt}\,({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ with generating functions $\Phi({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})+({\bf A}^{\prime}+{\bf a})\cdot{\mbox{\boldmath$\alpha$}}$:
${\bf A}={\bf A}^{\prime}+{\bf a}+{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}\Phi({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}}),\quad{\mbox{\boldmath$\alpha$}}^{\prime}={\mbox{\boldmath$\alpha$}}+{\mbox{\boldmath$\partial$}}_{{\bf A}^{\prime}}\Phi({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})$ (2.4)
with $\Phi$ chosen so that in the new coordinates $({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ the perturbation is ’weaker’; at the price that the new coordinates will cover a (much) smaller domain, inside ${\mathcal{C}}(\rho_{0},\kappa_{0})$.
@@
Suppose that a holomorphic $\Phi({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}})$ defines, via Eq.(2.4), in a domain $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})\in{\mathcal{C}}_{\widetilde{\rho},\widetilde{\kappa}}$, a symplectic map $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})\to({\bf A},{\mbox{\boldmath$\alpha$}})$ with image $\subset{\mathcal{C}}_{\rho_{0},\kappa_{0}}$.
Choose $\bf a$} as solution of the implicit equation ${\bf a}=-J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{\bf a}{\overline{f}}_{0}({\bf a})$ and insert Eq.(2.4) into $H_{0}$.
After easy&patient algebra, the Hamiltonian becomes,  444 $\alpha$ has to be imagined as a function of $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$, solution of the second implicit equation in Eq.2.4.  up to a constant, the sum of the four lines (that will be referred as $F_{j},j=0,\ldots,4$):
${(0):\quad}$ ${{}H^{\prime}({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})=\frac{1}{2}({\bf A}^{\prime}\cdot J_{0}{\bf A}^{\prime})+{\mbox{\boldmath$\omega$}}_{0}\cdot{\bf A}^{\prime}}$
${(1):\quad}$ ${{}+{\overline{f}}_{0}({\bf A}^{\prime}+{\bf a})-{\overline{f}}_{0}({\bf a})-{\mbox{\boldmath$\partial$}}_{{\bf a}}{\overline{f}}_{0}({\bf a})\cdot{\bf A}^{\prime}}$
${(2):\quad}$ ${{}+\Big{(}f_{0}({\bf A}^{\prime}+{\bf a}+{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}\Phi,{\mbox{\boldmath$\alpha$}})-f_{0}({\bf A}^{\prime}+{\bf a},{\mbox{\boldmath$\alpha$}})\Big{)}+\frac{1}{2}{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}\Phi\cdot J_{0}{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}\Phi}$
${(3):\quad}$ ${{}+({\mbox{\boldmath$\omega$}}_{0}+J_{0}({\bf A}^{\prime}+{\bf a}))\cdot{\mbox{\boldmath$\partial$}}_{{\mbox{\boldmath$\alpha$}}}\Phi+f_{0}({\bf A}^{\prime}+{\bf a},{\mbox{\boldmath$\alpha$}})-{\overline{f}}_{{\bf 0}}({\bf A}^{\prime}+{\bf a})}$
where a few terms have been added or subtracted (including free addition or subtraction of constants, like ${\mbox{\boldmath$\omega$}}_{0}\cdot{\bf a}$, $\frac{1}{2}J_{o}{\bf a}\cdot{\bf a}$ and ${\overline{f}}_{0}({\bf a})$): so that each line $F_{j},j>0$, can formally be estimated to have size $\sim O(\varepsilon_{0}^{2})$ in a smaller polydisk, while the line $(0)$ is the new unperturbed Hamiltonian (in the $({{\bf A}}^{\prime},{\mbox{\boldmath$\alpha$}}^{\prime})$ variables 4 ).  [^5] Heuristically: line (1) is a sum of quantities of order $\varepsilon_{0}$ but it is of second order in $\bf A$$^{\prime}$ and could be made of $O(\varepsilon_{0}^{2})$ by suitably restricting size of $\bf A$$^{\prime}$: remark that ${\overline{f}}_{0}({\bf a})$ is a constant added freely but the last term is there because of the choice ${\bf a}=-J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{{\bf a}}{\overline{f}}_{0}({\bf a})$. Line (2) can be made of $O(\varepsilon_{0}^{2})$ by choosing $\Phi$ of $O(\varepsilon_{0})$. Line (3) could even be set =0=0 and invites taking it a definition of $\Phi$ (if possible).
@@
${{}{\bf a}=-J_{0}^{-1}{\mbox{\boldmath$\partial$}}_{{\bf a}}{\overline{f}}_{0}({\bf a})}$
${{}\Phi({\bf A}^{\prime},{\mbox{\boldmath$\alpha$}})=-\sum_{{\bf 0}\neq{\mbox{\boldmath$\nu$}}\in\mathcal{Z}^{N}}\frac{f_{0,{\mbox{\boldmath$\nu$}}}({\bf A}^{\prime}+{\bf a})}{i({\mbox{\boldmath$\omega$}}_{0}\cdot{\mbox{\boldmath$\nu$}}+(({\bf A}^{\prime}+{\bf a})\cdot J_{0}{\mbox{\boldmath$\nu$}}))}e^{i{\mbox{\boldmath$\nu$}}\cdot{\mbox{\boldmath$\alpha$}}},}$
@@
  • In general an example of dimensional bound is a bound of a holomorphic function $f(\zeta)$ and of its derivatives at a point $\zeta$ in terms of its maximum $|f|$ in the definition domain and of the distance $d(\zeta)$ of $\zeta$ to the boundary: the $n^{th}$-derivative at $\zeta$ is bounded by $n!(\max|f|)\,d(\zeta)^{-n}$. Likewise a function $g(\zeta)$ holomorphic in an annulus bounded by the circles $\zeta=e^{i\alpha\pm\kappa},\alpha\in[0,2\pi]$, will have Fourier’s coefficients bounded by its maximum $\varepsilon$ in the annulus as $|g_{{\mbox{\boldmath$\nu$}}}|\leq\varepsilon\exp(-\kappa\,|{\mbox{\boldmath$\nu$}}|)$. The mentioned bounds will be called dimensional bounds.
  • In conclusion, if $\varepsilon=\max_{{\mathcal{C}}(\rho,\kappa)}|{\mbox{\boldmath$\partial$}}_{{\bf A}}F({{\bf A}},{\mbox{\boldmath$\alpha$}})|+\max_{{\mathcal{C}}(\rho,\kappa)}\frac{1}{\rho}|{\mbox{\boldmath$\partial$}}_{\mbox{\boldmath$\alpha$}}F({{\bf A}},{\mbox{\boldmath$\alpha$}})|$ and $F$ is holomorphic in ${\mathcal{C}}(\rho,\kappa)$, a dimensional bound on $F({{\bf A}},{\mbox{\boldmath$\alpha$}})$ for $({{\bf A}},{\mbox{\boldmath$\alpha$}})\in{\mathcal{C}}((1-\lambda)\rho,\kappa)$, $\lambda<1$, is, for ${\mbox{\boldmath$\nu$}}\neq 0$:
    $|{\mbox{\boldmath$\partial$}}_{{\bf A}}^{\bf n}{\mbox{\boldmath$\partial$}}_{{\bf A}}F({{\bf A}},{\mbox{\boldmath$\alpha$}})|<{\bf n}!\,(\lambda\rho)^{-|{\bf n}|}\,\varepsilon,\qquad|{\mbox{\boldmath$\nu$}}||F_{\mbox{\boldmath$\nu$}}({{\bf A}},{\mbox{\boldmath$\alpha$}})|<e^{-\kappa\,|{\mbox{\boldmath$\nu$}}|}\,\rho\,\varepsilon$ (2.7)
    where ${\bf n}\in\mathcal{Z}_{+}^{N},\,|{\bf n}|=\sum n_{i},\,{\bf n}!=\prod n_{i}!,\,{\mbox{\boldmath$\nu$}}\in\mathcal{Z}^{N},\,|{\mbox{\boldmath$\nu$}}|=\sum_{i}|\nu_{i}|$.  [^6] Remark that to estimate $\bf A$}}-derivatives the $\rho$ has to be reduced by a factor $1-\lambda$ while to estimate $\alpha$-derivatives $\kappa_{0}$ can be left unchanged, as such $m$-th $\alpha$-derivatives modify the second of Eq.2.7 by a factor $|{\mbox{\boldmath$\nu$}}|^{m}$.

### v2

@@ 1 Ergodicity
$\delta x(t)=x^{\prime}(\frac{i^{\prime}}{i}t)-x(t){\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}x^{\prime}(i^{\prime}\varphi)-x(i\varphi)\hfill$ (1.4)
@@
The variation $\delta({\overline{K}}+{\overline{V}}+{\overline{V}_{ext}}){\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}dU+dL$, with $U{\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}{\overline{K}}+{\overline{V}}$ and $dL{\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}d{\overline{V}_{ext}}$, has a natural interpretation of heat $\delta Q$ in/out in the process $x\to x^{\prime}$.  [^2] Because $\delta({\overline{K}}+{\overline{V}}){\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}dU$ and $\delta{\overline{V}_{ext}}{\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}\delta L$ are the variation of the internal energy $U$ and $dL$ is the work that the system performs: i.e. $\delta Q=\delta U+\delta L$.
@@
${{}={\bf-}2\frac{\delta i}{i}{\overline{K}}+{2\delta_{\xi}{{\overline{K}}}}+{\mbox{$\delta$}}_{\xi}{\bf({-{{\overline{K}}}}+{\overline{V}}+{\overline{V}_{ext}})}\equiv{\bf+}2\frac{\delta i}{i}{\overline{K}}+2\delta{{\overline{K}}}+{\bf 0}}$
@@
${\frac{\delta Q}{{\overline{K}}}}=+2\frac{\delta i}{i}+2\frac{\delta{\overline{K}}}{{\overline{K}}}={2\delta\log(i{\overline{K}})}{\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}{\bf\delta S}$ (1.7)
@@ 2 KAM
$H_{0}({\bf A},{\mbox{$\alpha$}})=\frac{1}{2}({\bf A}\cdot J_{0}{\bf A})+{\mbox{$\omega$}}_{0}\cdot{{\bf A}}+f_{0}({\bf A},{\mbox{$\alpha$}})\equiv h_{0}+f_{0}$ (2.1)
real analytic for $({\bf A},{\mbox{$\alpha$}})\in({\mathcal{D}}_{\rho}\times{\mathcal{T}}^{N})$, $N\geq 2$, with: ${\mathcal{D}}_{\rho}=\{{\bf A}\in R^{N},|A_{j}|<\rho\}$, ${\mathcal{T}}^{N}$ = $N$-dimensional torus $0,2\pi]^{N}$, identified with unit polydisk, $\{{\bf z}|\,z_{j}=e^{i\alpha_{j}},j=1,\ldots,N\}$, ${\mbox{$\omega$}}_{0}\in R^{N}$ and $J_{0}$ is a $N\times N$ matrix, positive for simplicity, with eigenvalues in $J_{0,-},J_{0,+}]$.
@@
${{\mathcal{C}}_{\rho_{0},\kappa_{0}}}$ ${{}{\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}\{({\bf A},{\bf z})||A_{j}|\,\leq\,\rho_{0},\ e^{-\kappa_{0}}\leq|e^{i\alpha_{j}}|\leq e^{\kappa_{0}},j=1,\ldots,N\}\subset{\mathcal{C}}^{2N}}$
@@
${||f}$ ${{}||_{\rho_{0},\kappa_{0}}{\mathrel{\mathop{\kern 0.0pt=}\limits^{def}}}\max_{{\mathcal{C}}_{\rho_{0},\kappa_{0}}}|f({\bf A},{\bf z})|,\qquad\forall f\ {\rm holomorphic\ in}\ {\mathcal{C}}_{\rho_{0},\kappa_{0}}}$
@@
The idea is to focus attention on the center of ${\mathcal{C}}_{\rho_{0},\kappa_{0}}$ where, if $\varepsilon_{0}=0$, a motion (“free motion”: $({\bf A},{\mbox{$\alpha$}})\to({\bf A},{\mbox{$\alpha$}}+{\mbox{$\omega$}}_{0}t)$) takes place which is quasi periodic “with spectrum” ${\mbox{$\omega$}}_{0}$. This is done by studying motions in a small polydisk ${\mathcal{C}}_{\widetilde{\rho},\widetilde{\kappa}}({\bf a})\subset{\mathcal{C}}_{\rho_{0},\kappa_{0}}$, eccentric if ${\bf a}\neq{\bf 0}$.
Thus motions starting in ${\mathcal{C}}(\widetilde{\rho},\widetilde{\kappa})({\bf a})\subset{\mathcal{C}}(\rho_{0},\kappa_{0})$ near the center of ${\mathcal{C}}(\rho_{0},\kappa_{0})$ can be studied as “through a microscope”: in the good cases (i.e. under suitable assumption on the initial parameters $J_{0},{\mbox{$\omega$}}_{0}$ and $f_{0}$) the Hamiltonian will turn out to be expressible (after a further coordinate change, to turn ${\mathcal{C}}(\widetilde{\rho},\widetilde{\kappa})({\bf a})$ into ${\mathcal{C}}(\rho_{1},\kappa_{1})$, centered at 0) in a form substantially closer to that of a quasi periodic oscillator (described by its “normal” Hamiltonian ${\mbox{$\omega$}}_{0}\cdot{\bf A}$, or ${\mbox{$\omega$}}_{0}\cdot{\bf A}+\frac{1}{2}{{\bf A}}\cdot J_{0}{{\bf A}}$, in the variables ${\bf A},{\mbox{$\alpha$}}$).
Iterating the process the Hamiltonian changes but, remaining analytic in a polydisk $C_{\rho_{n},\kappa_{n}}$, converges to that of a harmonic oscillator: the interpretation will be that, looking very carefully in the vicinity of the “unperturbed” torus ${\mathcal{T}}_{{\mbox{$\omega$}}_{0}}=\{{\bf A}={\bf 0},{\mbox{$\alpha$}}\in[0,2\pi]^{N}\}$, the perturbed Hamiltonian exhibits with increasing precision a harmonic motion with spectrum ${\mbox{$\omega$}}_{0}$.
@@
The “spectrum” ${\mbox{$\omega$}}_{0}$ will be supposed “diophantine”, i.e. for some $C>0$ it is, denoting $\mathcal{Z}^{N}$ the lattice of the integers, for all ${\bf 0}\neq{\mbox{$\nu$}}\in\mathcal{Z}^{N}$:
$|{\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}}|\,\geq\,C\,|{\mbox{$\nu$}}|^{-N},\qquad|{\mbox{$\nu$}}|\equiv\sum_{i=1}^{N}|\nu_{i}|>0$ (2.3)
The inequality Eq.2.3 will be repeatedly used to define canonical transformations $({{\bf A}},{\mbox{$\alpha$}})\,{\kern-1.0pt\leftarrow\kern-5.0pt\to\kern-1.0pt}\,({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})$ with generating functions $\Phi({\bf A}^{\prime},{\mbox{$\alpha$}})+({\bf A}^{\prime}+{\bf a})\cdot{\mbox{$\alpha$}}$:
${\bf A}={\bf A}^{\prime}+{\bf a}+{\mbox{$\partial$}}_{\mbox{$\alpha$}}\Phi({\bf A}^{\prime},{\mbox{$\alpha$}}),\quad{\mbox{$\alpha$}}^{\prime}={\mbox{$\alpha$}}+{\mbox{$\partial$}}_{{\bf A}^{\prime}}\Phi({\bf A}^{\prime},{\mbox{$\alpha$}})$ (2.4)
with $\Phi$ chosen so that in the new coordinates $({\bf A}^{\prime},{\mbox{$\alpha$}}^{\prime})$ the perturbation is ’weaker’; at the price that the new coordinates will cover a (much) smaller domain, inside ${\mathcal{C}}(\rho_{0},\kappa_{0})$.
@@
Suppose that a holomorphic $\Phi({{\bf A}}^{\prime},{\mbox{$\alpha$}})$ defines, via Eq.(2.4), in a domain $({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})\in{\mathcal{C}}_{\widetilde{\rho},\widetilde{\kappa}}$, a symplectic map $({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})\to({\bf A},{\mbox{$\alpha$}})$ with image $\subset{\mathcal{C}}_{\rho_{0},\kappa_{0}}$.
Choose $\bf a$} as solution of the implicit equation ${\bf a}=-J_{0}^{-1}{\mbox{$\partial$}}_{\bf a}{\overline{f}}_{0}({\bf a})$ and insert Eq.(2.4) into $H_{0}$.
[…]
${(0):\quad}$ ${{}H^{\prime}({\bf A}^{\prime},{\mbox{$\alpha$}}^{\prime})=\frac{1}{2}({\bf A}^{\prime}\cdot J_{0}{\bf A}^{\prime})+{\mbox{$\omega$}}_{0}\cdot{\bf A}^{\prime}}$
${(1):\quad}$ ${{}+{\overline{f}}_{0}({\bf A}^{\prime}+{\bf a})-{\overline{f}}_{0}({\bf a})-{\mbox{$\partial$}}_{{\bf a}}{\overline{f}}_{0}({\bf a})\cdot{\bf A}^{\prime}}$
${(2):\quad}$ ${{}+\Big(f_{0}({\bf A}^{\prime}+{\bf a}+{\mbox{$\partial$}}_{\mbox{$\alpha$}}\Phi,{\mbox{$\alpha$}})-f_{0}({\bf A}^{\prime}+{\bf a},{\mbox{$\alpha$}})\Big)+\frac{1}{2}{\mbox{$\partial$}}_{\mbox{$\alpha$}}\Phi\cdot J_{0}{\mbox{$\partial$}}_{\mbox{$\alpha$}}\Phi}$
${(3):\quad}$ ${{}+({\mbox{$\omega$}}_{0}+J_{0}({\bf A}^{\prime}+{\bf a}))\cdot{\mbox{$\partial$}}_{{\mbox{$\alpha$}}}\Phi+f_{0}({\bf A}^{\prime}+{\bf a},{\mbox{$\alpha$}})-{\overline{f}}_{{\bf 0}}({\bf A}^{\prime}+{\bf a})}$
where a few terms have been added or subtracted (including free addition or subtraction of constants, like ${\mbox{$\omega$}}_{0}\cdot{\bf a}$, $\frac{1}{2}J_{o}{\bf a}\cdot{\bf a}$ and ${\overline{f}}_{0}({\bf a})$): so that each line $F_{j},j>0$, can formally be estimated to have size $\sim O(\varepsilon_{0}^{2})$ in a smaller polydisk, while the line $(0)$ is the new unperturbed Hamiltonian (in the $({{\bf A}}^{\prime},{\mbox{$\alpha$}}^{\prime})$ variables 4 ).  [^5] Heuristically: line (1) is a sum of quantities of order $\varepsilon_{0}$ but it is of second order in $\bf A$$^{\prime}$ and could be made of $O(\varepsilon_{0}^{2})$ by suitably restricting size of $\bf A$$^{\prime}$: remark that ${\overline{f}}_{0}({\bf a})$ is a constant added freely but the last term is there because of the choice ${\bf a}=-J_{0}^{-1}{\mbox{$\partial$}}_{{\bf a}}{\overline{f}}_{0}({\bf a})$. Line (2) can be made of $O(\varepsilon_{0}^{2})$ by choosing $\Phi$ of $O(\varepsilon_{0})$. Line (3) could even be set =0=0 and invites taking it a definition of $\Phi$ (if possible).
@@
${{}{\bf a}=-J_{0}^{-1}{\mbox{$\partial$}}_{{\bf a}}{\overline{f}}_{0}({\bf a})}$
${{}\Phi({\bf A}^{\prime},{\mbox{$\alpha$}})=-\sum_{{\bf 0}\neq{\mbox{$\nu$}}\in\mathcal{Z}^{N}}\frac{f_{0,{\mbox{$\nu$}}}({\bf A}^{\prime}+{\bf a})}{i({\mbox{$\omega$}}_{0}\cdot{\mbox{$\nu$}}+(({\bf A}^{\prime}+{\bf a})\cdot J_{0}{\mbox{$\nu$}}))}e^{i{\mbox{$\nu$}}\cdot{\mbox{$\alpha$}}},}$
@@
  • In general an example of dimensional bound is a bound of a holomorphic function $f(\zeta)$ and of its derivatives at a point $\zeta$ in terms of its maximum $|f|$ in the definition domain and of the distance $d(\zeta)$ of $\zeta$ to the boundary: the $n^{th}$-derivative at $\zeta$ is bounded by $n!(\max|f|)\,d(\zeta)^{-n}$. Likewise a function $g(\zeta)$ holomorphic in an annulus bounded by the circles $\zeta=e^{i\alpha\pm\kappa},\alpha\in[0,2\pi]$, will have Fourier’s coefficients bounded by its maximum $\varepsilon$ in the annulus as $|g_{{\mbox{$\nu$}}}|\leq\varepsilon\exp(-\kappa\,|{\mbox{$\nu$}}|)$. The mentioned bounds will be called dimensional bounds.
  • In conclusion, if $\varepsilon=\max_{{\mathcal{C}}(\rho,\kappa)}|{\mbox{$\partial$}}_{{\bf A}}F({{\bf A}},{\mbox{$\alpha$}})|+\max_{{\mathcal{C}}(\rho,\kappa)}\frac{1}{\rho}|{\mbox{$\partial$}}_{\mbox{$\alpha$}}F({{\bf A}},{\mbox{$\alpha$}})|$ and $F$ is holomorphic in ${\mathcal{C}}(\rho,\kappa)$, a dimensional bound on $F({{\bf A}},{\mbox{$\alpha$}})$ for $({{\bf A}},{\mbox{$\alpha$}})\in{\mathcal{C}}((1-\lambda)\rho,\kappa)$, $\lambda<1$, is, for ${\mbox{$\nu$}}\neq 0$:
    $|{\mbox{$\partial$}}_{{\bf A}}^{\bf n}{\mbox{$\partial$}}_{{\bf A}}F({{\bf A}},{\mbox{$\alpha$}})|<{\bf n}!\,(\lambda\rho)^{-|{\bf n}|}\,\varepsilon,\qquad|{\mbox{$\nu$}}||F_{\mbox{$\nu$}}({{\bf A}},{\mbox{$\alpha$}})|<e^{-\kappa\,|{\mbox{$\nu$}}|}\,\rho\,\varepsilon$ (2.7)
    where ${\bf n}\in\mathcal{Z}_{+}^{N},\,|{\bf n}|=\sum n_{i},\,{\bf n}!=\prod n_{i}!,\,{\mbox{$\nu$}}\in\mathcal{Z}^{N},\,|{\mbox{$\nu$}}|=\sum_{i}|\nu_{i}|$.  [^6] Remark that to estimate $\bf A$}}-derivatives the $\rho$ has to be reduced by a factor $1-\lambda$ while to estimate $\alpha$-derivatives $\kappa_{0}$ can be left unchanged, as such $m$-th $\alpha$-derivatives modify the second of Eq.2.7 by a factor $|{\mbox{$\nu$}}|^{m}$.